Business and financial risk

The impact of financial leverage on return on equity and earnings per share

Consider the following case of Happy Turtle Transportation Company:

Suppose Happy Turtle Transportation Company is considering a project that will require $300,000 in assets.

The company is small, so it is exempt from the interest deduction limitation under the new tax law.
The project is expected to produce earnings before interest and taxes (EBIT) of $55,000.
Common equity outstanding will be 10,000 shares.
The company incurs a tax rate of 25%.

If the project is financed using 100% equity capital, then Happy Turtle Transportation Company’s return on equity (ROE) on the project will be12.38%   . In addition, Happy Turtle’s earnings per share (EPS) will be$3.10   .

Alternatively, Happy Turtle Transportation Company’s CFO is also considering financing the project with 50% debt and 50% equity capital. The interest rate on the company’s debt will be 13%. Because the company will finance only 50% of the project with equity, it will have only 5,000 shares outstanding. Happy Turtle Transportation Company’s ROE and the company’s EPS will be   if management decides to finance the project with 50% debt and 50% equity.

When a firm uses debt financing, the business risk exposure for the firm’s common shareholders willincrease   .
Answer #1

ROE = (EBIT – Interest)(1-Tax Rate)/Equity

When financed with 100% equity, ROE = 55,000(1-25%)/300,000 = 13.75%

EPS = (55,000)(1-25%)/10,000

= $4.125 per share

When financed with 50% Debt

ROE = (55,000 – 150,000*13%)(1-25%)/150,000

= 17.75%

EPS = 26,625/5,000

= 5.325 per share

INCREASE ROE